# SDP bounds on quantum codes

### v1

@@ 5.9. Self-dual stabilizer codes
One can give the following argument on the appearance of the Lovász number in the case of self-dual stabilizer codes. A stabilizer code corresponds to an abelian stabilizer group $S$ for which $-\mathds{1}\not\in S$. Given $s_{a},s_{b}\in S$, there exist $s_{c}\in S$ such that $s_{c}=s_{a}s_{b}$. A pure stabilizer code additionally satisfies $\operatorname{wt}(s_{i})\geq\delta$ for all $s_{i}\in S\setminus\mathds{1}$.
Recall that $\vartheta(G)$ upper bounds the independence number $\alpha(G)$, i.e., the maximum size of a set of disconnected vertices of a graph $G$. Let $G^{\prime}$ be the confusability graph of a self-dual code with the identity vertex removed. Then two vertices $a,b\in V$ are disconnected if they satisfy
(72) $E_{a}^{\dagger}E_{b}+E_{b}E_{a}^{\dagger}=0\quad\text{and}\quad\operatorname{wt}(E_{a}^{\dagger}),\operatorname{wt}(E_{b}),\operatorname{wt}(E_{a}^{\dagger}E_{b})\geq\delta\,.$
By setting $E_{c}=\pm E_{a}^{\dagger}E_{b}$, the equation above shows that the vertices $a,b,c\in V$ do not share any edges. Denote by $H\subseteq V$ a subset of disconnected vertices of the graph $G^{\prime}$ with maximum size $\alpha(G^{\prime})$. Any $a,b\in H$ then satisfies
(73) $E_{a}^{\dagger}E_{b}+E_{b}E_{a}^{\dagger}=0,\quad\operatorname{wt}(E_{a}^{\dagger}),\operatorname{wt}(E_{b})\geq\delta\quad\text{and}\quad c\in H\quad\text{if}\quad E_{c}=\pm E_{a}^{\dagger}E_{b}.$
The above equation allows us to state the following: Given a self-dual stabilizer code $S$, one can construct a disconnected subset $H_{S}\subseteq V$ with size $12^{n}-$ 1 by identifying each vertex $a\in V$ with a Pauli error $E_{a}$ such that $E_{a}=\pm s_{a}\in S\setminus\mathds{1}$. Conversely, given $H$, one can construct a subset of $\mathcal{E}_{n}$ by identifying each vertex $a\in V$ with $E_{a}$. Using Eq. (73), it can be seen that all elements of this set has weight equal or larger than $\delta$ and up to real phases, it forms an abelian group if one includes $+\mathds{1}$. But this is nothing else than a stabilizer group. Note that this stabilizer group can have at most $12^{n}-$ 1 nontrivial elements since the maximum number of Pauli errors which commutes between each other is $2^{n}$. Therefore, a self-dual stabilizer code exists, if and only if, $\alpha(G^{\prime})=2^{n}-1$. When $\alpha(G^{\prime})<2^{n}-1$, it is not possible to construct a stabilizer group with $2^{n}$ elements and therefore, a self-dual stabilizer code does not exist. Everything also holds for the subgraph $G^{\prime\prime}$ from Remark 11.
@@ 6.1. Automorphism groups
Let ${\boldsymbol{q}}=\{0,\dots,q-1\}$ be an alphabet and ${\boldsymbol{q}}^{n}$ be the set of strings of length $n$ formed from elements of $q$. Denote the Hamming distance between two elements $x,y\in{\boldsymbol{q}}^{n}$ by $d(x,y)$ the number of positions in which $x$ and $y$ differ. Then consider the set of automorphism $\operatorname{Aut}(q,n)$ of ${\boldsymbol{q}}^{n}$ that preserve the Hamming distance. It is known that $\operatorname{Aut}(q,n)$ consists of the set of permutations that permute the $n$ coordinates and that independently permute the alphabet at each coordinate [23]. Let $\operatorname{Aut}_{0}(q,n)$ be the subset of $\operatorname{Aut}(q,n)$ that leaves the all-zero string invariant. This consists of the set of permutations that permute the $n$ coordinates and independently permute the $q-1$ non-zero elements at each coordinate [23]. The group $\operatorname{Aut}_{0}(q,n)$ is isomorphic to the wreath product $S_{q-1}\wr S_{n}$.
The equivalence class of a classical code is generated by $\operatorname{Aut}(q,n)$. In turn, the equivalence class of a quantum code under local Clifford operations is generated by $\operatorname{Aut}_{0}(q,n)$. The reason is that the projector onto the code space has to remain positive-semidefinite under the equivalence operation, and consequentially the identity matrix has to be a fixed point.
@@ 6.2. Nonbinary Terwilliger algebra
Let $M_{i,j}^{t,p}$ a ${\boldsymbol{q}}^{n}\times{\boldsymbol{q}}^{n}$ matrix indexed by ${\boldsymbol{q}}^{n}$ and defined by
(75) $(M_{i,j}^{t,p})_{xy}=\begin{cases}1&\text{ if }\quad|s(x)|=i\,,\quad|s(x)|=j\,,\\ &\hskip 28.45274pt\lvert s(x)\cap s(y)\rvert=t\,,\quad\lvert\{m\,:\,x_{m}=y_{m}\neq 0\}\rvert=p\,,\vspace{0.2cm}\\ 0&\text{ otherwise}\,,\end{cases}$
@@
The range of integers, outside of which $M_{i,j}^{t,p}$ necessarily vanishes, is
@@
Let us give some motivation as to how the indexing with $(i,j,t,p)$ in Eq. (75) will be used later. Consider the three terms $E^{\dagger}$, $F$, and $EF^{\dagger}$ that appear in a single entry of the matrix $\Gamma$ [Eq. (56)]. Now suppose that $\operatorname{wt}(E)=i$ and $\operatorname{wt}(F)=j$. Then, set $t=|s(E)\cap s(F)|$ and $p=|\{m\,:\,(E)_{m}=(F)_{m}\neq\mathds{1}\}$ the number of tensor factors (coordinates) where both $E$ and $F$ have identical nontrivial Pauli elements. Then the overlap between the supports of $E$ and $F$ is $t$, and the number of places a nontrivial Pauli in $E$ cancels a nontrivial Pauli in $F$ is $p$. The quantity $t-p$ is then the number of places that a nontrivial Pauli in $E$ does not cancel a nontrivial Pauli in $F$. Note that if $E$ and $F$ commutes then $t-p$ is even and if $E$ and $F$ anticommutes then $t-p$ is odd. Finally, $\operatorname{wt}(EF^{\dagger})=i+j-t-p$.
@@ 6.3. Averaging over codewords
We now consider different averaging methods, whose results lie in the Terwilliger algebra. Recall that an automorphism $\pi\in\operatorname{Aut}(q,n)$ acts on a matrix $R$ by permuting its rows and columns, $\pi(R)_{xy}=R_{\pi(x)\pi(y)}$. Now let $\sigma_{v}$ be a distance-preserving automorphism that additionally satisfies $\sigma_{v}(v)=0$. With this, define $R^{v}$ as
@@
The effect of Eq. (79) is to average $R$ over all distance-preserving automorphisms that map codewords to the zero string.
@@
(82) $=\frac{1}{\left\lvert C\right\rvert}\left\langle\sum\limits_{v\in C}\sigma_{v}\!\left(R\right),M_{i,j}^{t,p}\right\rangle$
@@
(83) $\lambda^{t,p}_{i,j}=\frac{1}{\left\lvert C\right\rvert}\!\!\!\!\sum\limits_{\begin{subarray}{c}x,y\in{\boldsymbol{q}}^{n}\\ |s(x)|=i\\ |s(y)|=j\\ \lvert s(x)\cap s(y)\rvert=t\\ \lvert\{m\mid x_{m}=y_{m}\neq 0\}\rvert=p\\ \end{subarray}}\!\!\!\!\sum\limits_{v\in C}\sigma_{v}(R)_{xy}\,.$
since the only elements $\sigma_{v}(R)_{xb}(M_{i,j}^{t,p})_{xb}$ different to zero are the ones satisfying $|s(x)|=i$, $|s(x)|=j$, $\lvert s(x)\cap s(b)\rvert=t$ and |$\lvert\{m\mid x_{m}=b_{m}\neq 0\}\rvert=p$.
@@
Ref. [23, Proposition 9] defines $\langle\,\tilde{R},M_{i,j}^{t,p}\,\rangle=\frac{1}{|C|}\lambda^{t,p}_{i,j}$. However, we decide to define $\langle\,\tilde{R},M_{i,j}^{t,p}\,\rangle=\lambda^{t,p}_{i,j}$ since in the quantum case, $\tilde{\Gamma}$ is averaged over $\operatorname{Aut}_{0}(q,n)$ instead of averaging over $\operatorname{Aut}_{0}(q,n)$ and codewords (see Proposition  17).
@@
It can be shown that $\tilde{R}$ is positive semidefinite, if and only if
@@ 6.4. Averaging over the complement of a code
Since is invariant under all permutations $\sigma\in\operatorname{Aut}(q,n)$, the matrix $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is an element of the Bose–Mesner algebra [23, Eq. (40)]. Therefore, it can be written as
@@
Suppose that $\tilde{R}^{c}_{x0}=0$ for all $x$. This happens for example when $R=\chi\chi^{T}$ where $\chi$ is the indexing vector of a classical code. Ref. [23, Proposition 8] shows that
$\tilde{R}^{c}$ $=\frac{1}{q^{n}-|C|}(q^{n}\bar{R}-|C|\tilde{R})$
@@
Then, in analogy to Eq. (85), the $\tilde{R}^{c}\succeq 0$ is equivalent to
@@
6.5. The symmetry reduction of $\Gamma$
A natural mapping identifying $\mathds{1},Z,X,Y$ with $0,1,\omega,\omega^{2}\in\operatorname{\mathrm{GF}}(4)$ respectively maps, up to a phase, the multiplication of Pauli matrices to addition over $\operatorname{\mathrm{GF}}(4)$ [9]. That is, $E_{x+y}=i^{b}E_{x}E_{y}$ for some $b=0,1,2,3$. Therefore, we map $\mathcal{E}_{n}$ to $\operatorname{\mathrm{GF}}(4)^{n}$. In particular, $E_{0}=\mathds{1}^{\otimes n}$, where we write $E_{0}$ for $E_{0^{n}}$. Indexing the matrix $\Gamma$ from Eq. (56) with all elements from $\operatorname{\mathrm{GF}}(4)^{n}$ and eventually ${{\boldsymbol{q}}}^{n}$ with $q=4$, our aim is now to follow the averaging methods from Sections 6.3 and 6.4.
For any $v\in\operatorname{\mathrm{GF}}(4)^{n}$, define the automorphism $\sigma_{v}\in\operatorname{Aut}(\mathcal{E}_{n})$ acting on $\Gamma$ as
@@ 6.5. The symmetry reduction of $\Gamma$
(93) $={\langle E_{v}E_{x}^{\dagger}\rangle}{\langle E_{y}E_{v}^{\dagger}\rangle}{\langle E_{x}E_{y}^{\dagger}\rangle}\,.$
This maps the row and column indexed by $v$ to the one indexed by $0^{n}$ respectively. We can then average $\Gamma$ over a subset of $\operatorname{Aut}(\operatorname{\mathrm{GF}}(4),n)$, such that it can be expanded in the Terwilliger algebra. Note that averaging over distance-preserving automorphisms of $\operatorname{\mathrm{GF}}(4)^{n}$ has the same effect as over distance-preserving automorphisms of ${\boldsymbol{q}}^{n}$ with $q=4$, because the Hamming distance between code words is not affected by any group structure of the underlying alphabet. Thus with some abuse of notation, we define in analogy to Eq. (79) the matrix $\Gamma^{v}$ as
@@
where $Q$ is a subset from ${\boldsymbol{4}}^{n}$. The effect of Eq. (95) is to average $\Gamma$ over all distance-preserving automorphisms that map an element of $Q$ to $0^{n}$. Likewise, in analogy to Eq. (87) define the averaging over all elements from $\operatorname{Aut}(\mathcal{E}_{n})$ as
@@
We now consider averaging over different type of subsets $Q$. We first take $Q={\boldsymbol{4}}^{n}$. This is equivalent to averaging $\Gamma$ over all elements from $\operatorname{Aut}(4,n)$, which will lead to the quantum Delsarte bound.

### v2

@@ 5.9. Self-dual stabilizer codes
One can give the following argument on the appearance of the Lovász theta number in the case of self-dual stabilizer codes. Recall that a stabilizer code corresponds to an abelian subgroup of the $n$-qubit Pauli group for which $\mathds{1}\not\in S$. A pure stabilizer code additionally satisfies $\operatorname{wt}(s_{i})\geq\delta$ for all $s_{i}\in S\setminus\mathds{1}$.
Recall that $\vartheta(G)$ upper bounds the independence number $\alpha(G)$, i.e., the maximum size of a set with pairwise disconnected vertices. Let $G^{\prime}$ be the confusability graph of a self-dual code with the identity vertex removed. Then two vertices $a,b\in V$ are disconnected if and only if they satisfy all the following conditions,
(74) $E_{a}^{\dagger}E_{b}-E_{b}E_{a}^{\dagger}=0\quad\text{and}\quad\operatorname{wt}(E_{a}^{\dagger}),\operatorname{wt}(E_{b}),\operatorname{wt}(E_{a}^{\dagger}E_{b})\geq\delta\,.$
Given $s_{a},s_{b}\in S$, we also have $s_{c}=s_{a}s_{b}\in S$. Therefore, if $E_{c}=\pm E_{a}^{\dagger}E_{b}$, then $a,b,c\in V$ do not share any edges. Denote by $H\subseteq V$ a subset of disconnected vertices of the graph $G^{\prime}$ with maximum size $\alpha(G^{\prime})$. Any $a,b\in H$ then satisfies
(75) $E_{a}^{\dagger}E_{b}-E_{b}E_{a}^{\dagger}=0,\quad\operatorname{wt}(E_{a}^{\dagger}),\operatorname{wt}(E_{b})\geq\delta\,,\quad\text{and}\quad c\in H\quad\text{if}\quad E_{c}=\pm E_{a}^{\dagger}E_{b}.$
The above equation allows us to state the following: given a self-dual stabilizer code $S$, one can construct a disconnected subset $H_{S}\subseteq V$ with size $2^{n}-1$ by identifying each vertex $a\in V$ with a Pauli error $E_{a}$ such that $E_{a}=\pm s_{a}\in S\setminus\mathds{1}$. Conversely, given $H$, one can construct a subset of $\mathcal{E}_{n}$ by identifying each vertex $a\in V$ with $E_{a}$. Using Eq. (75), it can be seen that all elements of this set has weight equal or larger than $\delta$ and up to real phases, it forms an abelian group if one includes $\mathds{1}$. But this is nothing else than a stabilizer group. Note that this stabilizer group can have at most $2^{n}-1$ nontrivial elements since the maximum size of a commuting Pauli subgroup is $2^{n}$. Therefore, a self-dual stabilizer code exists if and only if $\alpha(G^{\prime})=2^{n}-1$. When $\alpha(G^{\prime})<2^{n}-1$, it is not possible to construct a stabilizer group with $2^{n}$ elements and therefore, a self-dual stabilizer code does not exist. The same reasoning applies to the subgraph $G^{\prime\prime}$ of Remark 12.
@@ 6.1. Automorphism groups
Let ${\boldsymbol{q}}=\{0,\dots,q-1\}$ be an alphabet and ${\boldsymbol{q}}^{n}$ be the set of strings of length $n$ formed from elements of $\boldsymbol{q}$}. Denote the Hamming distance between two elements $x,y\in{\boldsymbol{q}}^{n}$ by $d(x,y)$ the number of positions in which $x$ and $y$ differ. Then consider the set of automorphism $\operatorname{Aut}(q,n)$ of ${\boldsymbol{q}}^{n}$ that preserve the Hamming distance. It is known that $\operatorname{Aut}(q,n)$ consists of the set of permutations that permute the $n$ coordinates and that independently permute the alphabet at each coordinate [23].
Let $\operatorname{Aut}_{0}(q,n)$ be the subset of $\operatorname{Aut}(q,n)$ that leaves the all-zero string invariant. This consists of the set of permutations that permute the $n$ coordinates and independently permute the q1q-1 non-zero elements at each coordinate [23]. The group $\operatorname{Aut}_{0}(q,n)$ is isomorphic to the wreath product $S_{q-1}\wr S_{n}$.
The equivalence class of a classical code is generated by $\operatorname{Aut}(q,n)$. In turn, the equivalence class of a quantum code under local Clifford operations and permutations is generated by $\operatorname{Aut}_{0}(q,n)$  [^11] Usually equivalence under local unitary operations and permutations are considered, but due to our indexing of the moment matrix, the subgroup of local Clifford operations and permutations suffices.. The reason is that the projector onto the code space has to remain positive-semidefinite under the equivalence operation, and consequentially the identity matrix has to be a fixed point.
@@ 6.2. Nonbinary Terwilliger algebra
Let $M_{i,j}^{t,p}$ a $q^{n}\times q^{n}$ matrix indexed by ${\boldsymbol{q}}^{n}$ and defined by
[…]
@@
The range of tuples, outside of which $M_{i,j}^{t,p}$ necessarily vanishes, is
@@
Let us give some motivation as to how the indexing with $(i,j,t,p)$ in Eq. (77) will be used later. Consider the three terms $E^{\dagger}$, $F$, and $EF^{\dagger}$ that appear in a single entry of the matrix $\Gamma$ [Eq. (57)]. Now suppose that $\operatorname{wt}(E)=i$ and $\operatorname{wt}(F)=j$. Then, set $t=|s(E)\cap s(F)|$ and $p=|\{m\,:\,(E)_{m}=(F)_{m}\neq\mathds{1}\}$ the number of tensor factors (coordinates) where both $E$ and $F$ have identical nontrivial Pauli elements. Then the overlap between the supports of $E$ and $F$ is $t$, and the number of places a nontrivial Pauli in $E$ cancels a nontrivial Pauli in $F$ is $p$. The quantity tpt-p is then the number of places that a nontrivial Pauli in $E$ does not cancel a nontrivial Pauli in $F$. Note that if $E$ and $F$ commutes then tpt-p is even and if $E$ and $F$ anticommutes then tpt-p is odd. Finally, $\operatorname{wt}(EF^{\dagger})=i+j-t-p$.
@@ 6.3. Averaging over codewords
We now consider different averaging methods, whose results lie in the Terwilliger algebra. Recall that an automorphism $\pi\in\operatorname{Aut}(q,n)$ acts on a matrix $R$ by permuting its rows and columns, $\pi(R)_{xy}=R_{\pi^{-1}(x)\pi^{-1}(y)}$. Now for $v\in{\boldsymbol{q}}^{n}$, let $\sigma_{v}$ be a distance-preserving automorphism that additionally satisfies $\sigma_{v}(v)=0$. With this, define $R^{v}$ as
@@
The effect of Eq. (81) is to average $R$ over all distance-preserving automorphisms that map $v$ to the zero string.
@@
(84) $=\frac{1}{\left\lvert C\right\rvert}\left\langle\sum\limits_{v\in C}\sigma_{v}\!\left(R\right),M_{i,j}^{t,p}\right\rangle\,.$
@@
(85) $\lambda^{t,p}_{i,j}=\frac{1}{\left\lvert C\right\rvert}\!\!\!\!\sum\limits_{\begin{subarray}{c}x,y\in{\boldsymbol{q}}^{n}\\ |s(x)|=i\\ |s(y)|=j\\ \lvert s(x)\cap s(y)\rvert=t\\ \lvert\{m\mid x_{m}=y_{m}\neq 0\}\rvert=p\\ \end{subarray}}\!\!\!\!\sum\limits_{v\in C}\sigma_{v}(R)_{xy}\,$,
since the only elements $\sigma_{v}(R)_{xy}(M_{i,j}^{t,p})_{xy}$ not equal to zero are the ones satisfying $|s(x)|=i$, $|s(x)|=j$, $\lvert s(x)\cap s(y)\rvert=t$ and $\lvert\{m\mid x_{m}=y_{m}\neq 0\}\rvert=p$.
@@
Ref. [23, Proposition 9] defines $\langle\,\tilde{R},M_{i,j}^{t,p}\,\rangle=\frac{1}{|C|}\lambda^{t,p}_{i,j}$. However, we decide to define $\langle\,\tilde{R},M_{i,j}^{t,p}\,\rangle=\lambda^{t,p}_{i,j}$ since in the quantum case, $\tilde{\Gamma}$ is averaged over $\operatorname{Aut}_{0}(q,n)$ instead of averaging over both $\operatorname{Aut}_{0}(q,n)$ and codewords (see Proposition  18).
@@
It can be shown that $\tilde{R}$ is positive semidefinite if and only if
@@ 6.4. Averaging over the complement of a code
Since $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is invariant under all permutations $\sigma\in\operatorname{Aut}(q,n)$, the matrix $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is an element of the Bose–Mesner algebra [23, Eq. (40)]. Therefore, it can be written as
@@
Suppose that $\tilde{R}^{c}_{x0}=0$ for all $x$. This happens for example when $R=\chi\chi^{T}$ where $\chi$ is the incidence vector of a classical code. Ref. [23, Proposition 8] shows that
[…]
@@
Then, in analogy to Eq. (87), $\tilde{R}^{c}\succeq 0$ is equivalent to
@@
6.5. Symmetry reduction of $\Gamma$
A natural mapping identifying the matrices $I,Z,X,Y$ with $0,1,\alpha,\alpha^{2}\in\operatorname{\mathrm{GF}}(4)$ respectively, maps, up to a phase, the multiplication of Pauli matrices to addition over $\operatorname{\mathrm{GF}}(4)$ [9]. That is, $E_{x+y}=\omega_{xy}E_{x}E_{y}$ with $\omega_{xy}\in\{\pm i,\pm 1\}$, and where $E_{-x}=E_{x}^{\dagger}$. Therefore, we map $\mathcal{E}_{n}$ to $\operatorname{\mathrm{GF}}(4)^{n}$. In particular, $E_{0}=I^{\otimes n}$, where we write $E_{0}$ for $E_{0^{n}}$. Indexing the matrix $\Gamma$ from Eq. (57) with all elements from $\operatorname{\mathrm{GF}}(4)^{n}$ (and eventually ${{\boldsymbol{q}}}^{n}$ with $q=4$), our aim is now to follow the averaging methods from Sections 6.3 and 6.4.
For any $v\in\operatorname{\mathrm{GF}}(4)^{n}$, define $\sigma_{v}$ to be a distance-preserving automorphism on $\operatorname{\mathrm{GF}}(4)^{n}$. It acts on $\Gamma$ as
@@ 6.5. Symmetry reduction of $\Gamma$
(95) $={\langle E_{v}E_{x}^{\dagger}\rangle}{\langle E_{y}E_{v}^{\dagger}\rangle}{\langle E_{x}E_{y}^{\dagger}\rangle}\,$,
where the third equation follows from $\bar{\omega}_{x(-v)}\omega_{y(-v)}\omega_{x(-v)}\bar{\omega}_{y(-v)}=1$. This maps the row and column indexed by $v$ to the one indexed by $0^{n}$ respectively. We can then average $\Gamma$ over a subset of distance-preserving automorphisms on $\operatorname{\mathrm{GF}}(4)^{n}$ such that it can be expanded in the Terwilliger algebra.
Note that the averaging of $\Gamma$ over distance-preserving automorphisms of $\operatorname{\mathrm{GF}}(4)^{n}$ has the same effect as averaging over distance-preserving automorphisms of ${\boldsymbol{q}}^{n}$ with $q=4$. This is because the weight of $E_{x}^{\dagger}E_{y}$ (corresponding to the distance between $x$ and $y$) is not affected by any group structure of the underlying alphabet. More formally, the weight is characterized by $\operatorname{wt}(E_{x}E_{y}^{\dagger})=i+j-t-p$ with $i,j,t,p$ defined in Remark 13, which only depends on the alphabet and not the group structure. Thus, we can map $\operatorname{\mathrm{GF}}(4)^{n}$ to ${\boldsymbol{4}}^{n}$.
With some abuse of notation, we define in analogy to Eq. (81) the matrix $\Gamma^{v}$ as
@@
where $Q$ is a subset of ${\boldsymbol{4}}^{n}$. The effect of Eq. (97) is to average $\Gamma$ over all distance-preserving automorphisms that map an element of $Q$ to $0^{n}$. Likewise, in analogy to Eq. (89) define the averaging over all elements from $\operatorname{Aut}(4,n)$ as
@@
We now consider averaging over different subsets $Q$. We first take $Q={\boldsymbol{4}}^{n}$. This is equivalent to averaging $\Gamma$ over all elements from $\operatorname{Aut}(4,n)$, which will lead to the quantum Delsarte bound.